OpenStudy (anonymous):

Let y=2sqrtx . Find the change in y, deltay when x=5 and deltax=0.4 Find the differential dy when x=5 and dx=0.4

OpenStudy (ash2326):

\[y=2\sqrt x\] Can you find \(\frac{dy}{dx}\)?

OpenStudy (anonymous):

yeah i got x^-1/2dx

OpenStudy (ash2326):

good, so \[\large \frac{dy}{dx}=(x)^{-\frac 12}\] Find the change in y when "deltay when x=5 and deltax=0.4" so \[dx=0.4\] \[x=5\] find value of dy? just plugin the no.s

OpenStudy (agent0smith):

Do the same thing as last time :) just don't forget the dx :P

OpenStudy (anonymous):

i got dy=0.178885 is that correct

OpenStudy (agent0smith):

That's right.

OpenStudy (ash2326):

you're right

OpenStudy (agent0smith):

dy when x=5 and dx=0.4 i'd guess this one is meant to be 0.04 again. So just plug in 0.04 as dx this time.

OpenStudy (anonymous):

no i think we have to do something else cause its asking for the change in y i wanted to use an equation i saw in class its f(x+delta x)-f(x)

OpenStudy (agent0smith):

The change in y *is* the delta y. Delta means change.

OpenStudy (agent0smith):

Oh i think i see what you mean, though.

OpenStudy (ash2326):

\[\frac{f(x+\Delta x)-f(x)}{(x+\Delta x-x)}=f'(x) \]

OpenStudy (anonymous):

yeah i just used that equation and i got it right thanks you i just had to find i guess is the distance that x had change

OpenStudy (agent0smith):

Oh yeah i think you wanted the amount y changed to, ie the initial and final y. Not the actual dy.

OpenStudy (anonymous):

yeah thats what it is well thanks!! :)

OpenStudy (agent0smith):

For that, yes you just add dy to whatever y is when x=5.